The urn contains 2 white and 2 black balls.A ball is drawn at random.If it is white,it is not replaced into urn.Otherwise,it is replaced with another ball of the same colour.The process is repeated.Find probability that third ball drawn is black

Grade:12th pass

1 Answers

yaswanth
9 years ago
First we have to check all possible cases.for the third ball to be black,first ball can be 
white or black.second ball can be white or black.
case..1:first white second white third black
             2c1.1c1.2c1=4
case:2 :first white second black third black
             2c1.2c1.2c1=8
case 3:first black second white third black
            2c1.2c1.2c1=8
case 4:first black second black third black
            2c1.2c1.2c1=8
Favourable cases =4+8+8+8=28
    Probability =Favourable cases/total cases=28/44=7/11
if we have to add another black if black is drawn then answer is different
but it follows same procedure.actually the question doesn't have clarity in this case